# Slamming: Training a Speech Language Model on One GPU in a Day

### v1

@@ 5 Final Recipe
Building on these empirical findings, we develop the final Slam recipe. Using it, we train SLMs based on Qwen2.52.52.52.5-$0.5$B. We then compare Slam to the TWIST model family across various sizes: $350$M, $1.3$B, $7$B, and $13$B. We also present results for TWIST-$350$M using our computational constraints but following TWIST’s original training recipe, along with our synthetic data. Finally, we report results for the top-performing model from  Cuervo and Marxer (2024), including their predicted optimal performance under our compute budget based on SLM scaling laws. Results are reported in Table 2. The results indicate that Slam delivers performance that is either superior or on par with baseline models while requiring significantly fewer computational resources (e.g., a single A5000500050005000 for a day compared to $160$ days on a V100100100100).
@@ 6 Increasing Compute
Similarly to Geiping and Goldstein (2023), we analyse whether the proposed approach holds well also in increased compute budget. We opt for $48$ hours on $2$ A100100100100 GPUs as a reasonable academic budget for larger scale tests, and represents $\sim 10$ times more compute than the Slamming setting. We use exactly the same Slam recipe for more steps, and increase the batch size times $2$. We provide the full results in Table 3. We note that the performance continues to improve across all metrics, also outperforming methods which have far larger compute scales. We note that DPO training on synthetic data for $2$ epochs, notably boosts performance.
7 Conclusion
In this work we show that training high quality SLMs with a very modest compute budget, is feasible. We give these main guidelines: (i) Do not skimp on the model - not all model families are born equal and the TWIST initialisation exaggerates this, thus it is worth selecting a stronger / bigger text-LM even if it means less tokens. we found Qwen2.52.52.52.5 to be a good choice; (ii) Utilise synthetic training data - pre-training on data generated with TTS helps a lot; (iii) Go beyond next token prediction - we found that DPO boosts performance notably even when using synthetic data, and as little as $30$ minutes training massively improves results; (iv) Optimise hyper-parameters - as researchers we often dis-regard this stage, yet we found that tuning learning rate schedulers and optimising code efficiency can improve results notably. We hope that these insights, and open source resources will be of use to the research community in furthering research into remaining open questions in SLMs.
Limitations
@@ Appendix A Full Slam Recipe
We provide below the full training recipe, including hyperparameters for the best, Slam recipe. In Table  4 we see the Slam (-DPO) pre-training recipe and in Table  5 we see the Slam DPO training recipe.
Table 4: Slam (-DPO) Pre Training Recipe
@@
Data Librilight + Librispeech + sTinyStories
@@
Learning Rate Scheduler cosine with min $55e-$ 5
@@
Table 5: Slam DPO Training Recipe
@@
Data SpokenSwag with auto-bleu smaller than 0.3
@@ Appendix B Model Sizes
Model Name Number of Params
@@ Appendix C Dataset Statistics
We also investigate diverse datasets for pre-training: SWC Baumann et al. (2018), Tedlium Hernandez et al. (2018), PeopleSpeech Galvez et al. (2021) and VoxPopuli Wang et al. (2021b). We only take English subsets for all datasets, yet they can still contain diverse accents. These datasets are in the following domains SWC - read Wikipedia articles, Tedlium - short lectures, PeopleSpeech - diverse data including many local council gatherings etc, VoxPopuli - from European Parliament meetings. For SWC specifically, we use the text alignment to create chunks, remove silence from the audio and remove mis-aligned chunks. We use full training splits where provided, otherwise splitting 99% for training. The dataset sizes are described in 7.
Table 7: Dataset train set sizes that we use.
Dataset Number of Hours Number of Tokens
@@
Appendix D AI Tool Usage

### v2

@@ 5 Final Recipe
Building on these empirical findings, we develop the final Slam recipe. Using it, we train SLMs based on Qwen2.52.52.52.5-$0.5$B. We then compare Slam to the TWIST model family across various sizes: $350$M, $1.3$B, $7$B, and $13$B. We also present results for TWIST-$350$M using our computational constraints but following TWIST’s original training recipe, along with our synthetic data. Finally, we report results for the top-performing model from  Cuervo and Marxer (2024), including their predicted optimal performance under our compute budget based on SLM scaling laws. Results are reported in Table 2. The results indicate that Slam delivers performance that is either superior or on par with baseline models while requiring significantly fewer computational resources (e.g., a single A5000500050005000 for a day compared to $160$ days on a V100100100100). Transcribed generated examples by Slam can be seen in Appendix D.4.
To show that the Slam models do not overfit a single domain (audiobooks/stories), we provide results for GenPPL on a different domain. This can be seen in Appendix  D.3.
We further evaluate the quality of the generated audio using Mosnet Cooper et al. (2022), similarly to Align-SLM. Results are presented in Appendix D.2. As the quality of the generated audio is mainly affected by the vocoder, which is identical across evaluated methods, results are comparable. Interestingly, TWIST $1.3$B and TWIST $7$B achieve slightly worse scores.
@@ 6 Increasing Compute
Similarly to Geiping and Goldstein (2023), we analyse whether the proposed approach holds well also in increased compute budget. We opt for $48$ hours on $2$ A100100100100 GPUs as a reasonable academic budget for larger scale tests, and represents $\sim 10$ times more compute than the Slamming setting. We use exactly the same Slam recipe for more steps, and increase the batch size times $2$. We provide the full results in Table 3. We note that the performance continues to improve across all metrics, also outperforming methods which have far larger compute scales. We note that DPO training on synthetic data for $2$ epochs, notably boosts performance. Transcribed generated examples by Slam (scaled) can be seen in Appendix D.4
We also wish to assess whether our suggested recipe holds for larger models, thus we evaluate training a larger Qwen2.52.52.52.5 text LM as the base model. We use Qwen2.51.52.51.52.5-1.52.5 - 1.5B for the same compute budget as above - i.e two A100100100100 GPUs for $48$ hours. All training details are identical, but of course the larger model was trained for less steps (and tokens). We provide results from this model, denoted Slam (large) in Table 3. Results show that this model even outperforms the smaller model for the same compute budget. This demonstrates that the Slam recipe holds for larger models, and re-iterates the importance of quality models even at the expense of less training tokens for this setup.
7 Limitations
@@
8 Conclusion
In this work we show that training high quality SLMs with a very modest compute budget, is feasible. We give these main guidelines:
  1. 1.
    Do not skimp on the model - not all model families are born equal and the TWIST initialisation exaggerates this, thus it is worth selecting a stronger / bigger text-LM even if it means less tokens. we found Qwen2.52.52.52.5 to be a good choice.
  2. 2.
    Utilise synthetic training data - pre-training on data generated with TTS helps a lot.
  3. 3.
    Go beyond next token prediction - we found that DPO boosts performance notably even when using synthetic data, and as little as $30$ minutes training massively improves results.
  4. 4.
    Optimise hyper-parameters - as researchers we often dis-regard this stage, yet we found that tuning learning rate schedulers and optimising code efficiency can improve results notably.
We hope that these insights, and open source resources will be of use to the community in furthering SLM research.
@@ 8 Conclusion
Acknowledgements.
This research work was supported by ISF grant 2049/22.
@@ Appendix A Full Slam Recipe
We provide below the full training recipe, including hyperparameters for the best, Slam recipe. In Table 4 we see the Slam (-DPO) pre-training recipe and in Table 5 we see the Slam DPO training recipe. Table 6 provides the sampling hyper-parameters used for calculating the generative metrics. Note that some of the generated samples in the demo page were created with a higher maximum token limit.
Table 4: Slam (-DPO) pre-training recipe.
@@
Data Librilight, Librispeech, sTinyStories
@@
LR Scheduler cosine with min $55e-$ 5
@@
Table 5: Slam DPO training recipe.
@@
Data SpokenSwag with auto-bleu0.3absent0.3$\leq$ 0.3≤ 0.3
@@
Table 6: Slam sampling parameters.
Parameter Value
Temperature $0.8$
Top-K $25$
Max New Tokens $150$
Repetition Penalty $1.1$
@@ Appendix B Model Sizes
Model Name Params
@@ Appendix C Dataset Statistics
We also investigate diverse datasets for pre-training: SWC Baumann et al. (2018), Tedlium Hernandez et al. (2018), PeopleSpeech Galvez et al. (2021) and VoxPopuli Wang et al. (2021b). We only take English subsets for all datasets, yet they can still contain diverse accents. These datasets are in the following domains SWC - read Wikipedia articles, Tedlium - short lectures, PeopleSpeech - diverse data including many local council gatherings etc, VoxPopuli - from European Parliament meetings. For SWC specifically, we use the text alignment to create chunks, remove silence from the audio and remove mis-aligned chunks. We use full training splits where provided, otherwise splitting 99% for training. The dataset sizes are described in Table 8.
Table 8: Training set size for used datasets.
Dataset Hours Tokens
@@
Appendix D Additional Results
D.1 Context Length and Batch Size Ablation
In Table 9 we see results for ablations of context length and effective batch size
Table 9: Performance on sBlimp, tStoryCloze (tSC), sStoryCloze (sSC) and validation loss across different context lengths and effective batch sizes. using Qwen2.5-0.5B
[…]
512 128 56.13 76.91 55.53 1.67
512 256 56.56 77.49 56.33 1.66
1024 256 56.43 78.88 55.69 1.65
1024 128 56.45 78.01 55.59 1.64
D.2 MOS Proxy Results
For completeness we also provide MOS proxy results for our models compared to TWIST and Align-SLM models. We follow a similar setup to Lin et al. (2024) and use MOSnet to test the audio’s generation quality of our models. It is important to note that we use the same vocoder as TWIST and Align-SLM. The results can be seen in Table 10.
Table 10: MOSnet scores for various models.
Model MOSnet $\uparrow$
Ground Truth 4.28
TWIST-350M (Hassid et al., 2024) 4.07
TWIST-1.3B (Hassid et al., 2024) 3.83
TWIST-7B (Hassid et al., 2024) 3.85
Align-SLM-1.3B (Lin et al., 2024) 4.05
Align-SLM-7B (Lin et al., 2024) 4.09
Slam 4.11
Slam (scaled) 4.07
Slam (large) 4.05
D.3 GenPPL on Different Domain
In order to evaluate the generalisability of our approach to diverse domains, we calculate GenPPL for a dataset from a different domain. We compare our results to TWIST of various sizes, which were trained on this exact dataset (perhaps even overlapping samples as no official training set was published). We use the same setup for GenPPL as described in section 3, but we use People Speech test set Galvez et al. (2021) as prompts. Results in Table 11, show that Slam performs comparably or better to TWIST models of larger scale and more train compute, despite the fact they were explicitly trained on this dataset. These results highlight the efficacy of the Slam recipe beyond a single domain.
Table 11: GenPPL results on the People Speech Dataset.
Model GenPPL $\downarrow$ BLEU $\downarrow$
TWIST-350M (Hassid et al., 2024) 162.89 2.36
TWIST-1.3B (Hassid et al., 2024) 155.51 2.86
TWIST-7B (Hassid et al., 2024) 128.74 2.72
Slam 171.84 4.89
Slam (scaled) 128.78 3.99
Slam (large) 117.49 3.72
D.4 Text Generation Examples
For completeness we provide some transcriptions for some of the generated examples (generated by Slam and Slam (Scaled)). The prompts and the transcriptions of the generated samples can be seen in Table 12.
Table 12: transcriptions of generated continuations from Slam and Slam Scaled for various prompts.
Prompt Slam Slam Scaled
Johnny found an old coin on the subway floor… and decided to bring it home. Then he came up by a different subway floor, which opened into a courtyard. Johnny picked up the coin and threw it at the doorstep of the courtyard. He picked it up and looked at the size and shape of the coin. It must be valuable, remembering his promise to Johnny, had it put very carefully in his pocket. But now he felt as if somebody had given him a special coin in return.
I am a real boy he said He looked at Sarah Weinfeild and smiled. It was quite impossible for Sarah to look at him any longer, and had no suspicions about his feelings for a moment. So off he went to school in the spring time, sitting on the soft grass near the trees where the rain fell in torrents. As the sun began to set, the birds sang their lovely songs and the children clapped their hands with Joe and Jack Pensilvenya.
I had a great day today playing I saw a little boy playing too. He said hello to me, I went to say hello to you. I said hello back. I thought he was very nice, I asked him if he wanted to play with me? He said yes, and they played together telling jokes and laughter in the quiet little world. They had a great time together, and when it was time to go home I said bye to the little boy and went to sleep. end. I love playing the flute, and almost forgot my troubles in the warm sunshine as I set the flute for miles around in the air practicing my favorite songs. My flute singing out in answer to the beautiful sounds from the world around me. At length, the song ended and I stood up and clapped my hands, then we passed on through the forest, over the meadow and along the river bank until we came to a beautiful valley.
Appendix E AI Tool Usage